OpenStudy (anonymous):

for f(x) = 1/(x^2-3), find f(2+h)

OpenStudy (michele_laino):

we have to replace \(x=2+h\), so we get: \[f\left( {2 + h} \right) = \frac{1}{{{{\left( {2 + h} \right)}^2} - 3}} = ...?\]

OpenStudy (anonymous):

I'm pretty lost on how to do this, I have my equation set up just like that, but the )2+h) part throws me off so much

OpenStudy (anonymous):

(2+h)**

OpenStudy (michele_laino):

hint: if we develop the square of binomial, we can write this: \[\begin{gathered} f\left( {2 + h} \right) = \frac{1}{{{{\left( {2 + h} \right)}^2} - 3}} = \hfill \\ \hfill \\ = \frac{1}{{4 + {h^2} - 4h - 3}} = ...? \hfill \\ \end{gathered} \]

OpenStudy (er.mohd.amir):

(2+h)^2=4+h^2+4h

OpenStudy (michele_laino):

oops.. I have made a typo: \[\begin{gathered} f\left( {2 + h} \right) = \frac{1}{{{{\left( {2 + h} \right)}^2} - 3}} = \hfill \\ \hfill \\ = \frac{1}{{4 + {h^2} + 4h - 3}} = ...? \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

Should I have ended up with \[1\div h^2+4h-3\]?

OpenStudy (anonymous):

I feel like I'm doing this very wrong

OpenStudy (michele_laino):

no, please you have to simplify similar terms, so we have to compute: 4-3=...?

OpenStudy (anonymous):

4-3=1

OpenStudy (michele_laino):

correct! so, here is the final answer: \[\begin{gathered} f\left( {2 + h} \right) = \frac{1}{{{{\left( {2 + h} \right)}^2} - 3}} = \hfill \\ \hfill \\ = \frac{1}{{4 + {h^2} + 4h - 3}} = \hfill \\ \hfill \\ = \frac{1}{{{h^2} + 4h + 1}} \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

Thank you!!!

OpenStudy (michele_laino):

:)